# Evaluating Similitude and Robustness of Deep Image Denoising Models via Adversarial Attack

### v1

@@ IV-C Color image denoising adversarial attack
Under the constraint of denoising-PGD, 3-channel color RGB adversarial samples are generated for plug-and-play model DPIR with noise level set to 25, maximum perturbation $\varepsilon$ set to 0.012, and 5 iterative steps.
Figure 9: Effect of denoising-PGD for color RGB images (butterfly).
@@
As can be seen in Fig. 9, the addition of adversarial noise does not make a perceptible difference to the noise image, but makes the model denoised substantially less effective. We can see that PSNR decreases by 1.1220 dB, MAE increases by 0.5151, and SSIM decreases by about 0.0082. Visually, the enlarged areas in the denoised image produce the wrong texture, and many image details are erased.
@@
In the experiments to test the transferability, we selected multiple types of high-span deep image denoising models for color RGB image denoising: non-blind denoising methods FFDNet and BRDNet, blind denoising methods Noise2Noise and FAN, plug-and-play methods DPIR and CurvPnP. The adversarial samples generated for DPIR under denoising-PGD are tested on the above models.
@@
As can be seen from Table 5, the transferability of denoising-PGD on RGB images generated for DPIR is less than the transferability on grayscale images generated for DnCNN. The adversarial samples show a higher adversarial effect on the same type of denoising model, with PSNR decreasing by 0.5051 dB, MAE increasing by 0.4069, and SSIM decreasing by 0.0073 for CurvPnP. While for other kinds of adversarial samples, the adversarial effect decreases, but are still transferable. Based on this, we can say that the best transferable adversarial samples can be obtained by applying the image denoising adversarial attack method to a non-blind deep image denoising model. The observation also coincides with the result in section III.C, i.e., the non-blind denoising model has the smallest adversarial space.
IV-D Deep image denoising adversarial training
@@
We used denoisng-PGD at an iteration number of 5 and $\varepsilon=0.012$ to generate the adversarial noise images. Mix the adversarial noise images and Gaussian noise images in 1:1 ratio and feed into the model training phase. The adversarial samples re-generated by denoising-PGD on DnCNN and DnCNN-B are integrated as a test set for testing the resistance of the model after adversarial training against adversarial samples. As can be seen in Fig. 10, the PSNR decremental are reduced from the original 0.8862 dB to 0.1609 dB, and visually denoised adversarial samples become closer to original images. Evaluation of the effect of adversarial training is showed in Table 6. The difference in the average denoising effect before and after adversarial training is reduced from 0.807 dB to an impressive 0.117 dB. Surprisingly, the adversarial trained denoising model doesn’t show any performance decreasing on original Gaussian noise images, as the main drawback of adversarial training is performance decreasing on the original task. Fig. 11 shows that the adversarial trained model has even better denoising effect on pure Gaussian noise images than vanilla deep image denoising model. The adversarial trained model can reduce artifacts which are commonly appeared in deep image denoising models.
Figure 11: Denoising effect of the adversarial training model and the original model when faced with the same Gaussian noise image.
@@ V-A Image denoising adversarial perturbation on clean images
Figure 12: Demonstration of the effect of generating adversarial samples based on the original and noise images. (a) Based on the clean image (b) Based on the noise image.
@@
In summary, the image denoising network do not have adversarial in the case that the input clean image generates adversarial instances. We conjecture that similar to the classification boundary of image classification task, image denoising also has a denoising space, and the outide of the space is the part of the network that is difficult to denoise. Therefore, in this paper, we use the noise map with Gaussian noise added to the input to generate the adversarial samples.
V-B Image denoising adversarial attack in the classic method
@@
Figure 13: Effectiveness of BM3D in the face of image denoising adversarial attacks, where the adversarial samples are generated by the DnCNN network under denoising-PGD. The test values are PSNR/MAE/SSIM from left to right, respectively.
@@ V-C L2-denoising-PGD image denoising adversarial attack
  INPUT:
       A gauss noise example $x$ with clean example $y$; a denoiser $D$ with loss function $J$; perturbation size $\varepsilon$; maximum iterations $T$.
  OUTPUT:
       An adversarial sample ${x}^{\prime}$.
@@
V-D The effect of L2-denosing-PGD adversarial attack
In order to eliminate the disturbing pitfall that the random perturbation of the PGD method may have an impact on the noise distribution, the L2-denosing-PGD image denoising adversarial attack method is proposed. As can be seen from the distribution comparison plot in Fig. 14, the adversarial noise of denoising-PGD still has some influence on the distribution, which overlaps more with the Gaussian distribution curve after the L2 constraint is applied. In other words, the noise distribution of denoising-PGD is close to the Gaussian distribution assumption, while the noise distribution of the L2-denoising-PGD is more consistent with the Gaussian distribution assumption after the constraint is applied.
Figure 14: Visual comparison of the noise distribution of the adversarial samples generated under denoising-PGD with the Gaussian noise corresponding to the noise level. (Upper)denoising-PGD, W-distance=0.0024. (Lower)L2-denoising-PGD, W-distance=0.0017.
Figure 15: The effect diagram of the image denoising adversarial attack method is shown. (a)denoising-PGD. (b)L2-denoising-PGD.
The L2-denosing-PGD image denoising adversarial attack method also attaches an L2 constraint to the perturbation range of the noise, which further enhances the imperceptibility of the adversarial perturbation and further reduces the impact on the original image. As can be seen from Fig. 15, the effect of the adversarial under such a constraint that conforms to the Gaussian distribution noise assumption and limits the noise perturbation range is very little affected, and the psnr residual is 0.8051 dB, which is 0.2584 dB lower than the case without the constraint, and the image still has a significant adversarial effect.
In addition, by using the L2-denosing-PGD on different kinds of deep neural denoising networks to generate grayscale adversarial samples and color adversarial samples and testing their adversarial and transferability on the remaining classes of deep neural denoising networks, we can witness that the method is still superior in maintaining the noise distribution, eliminating the concern that the adversarial effect is significantly related to the change of the noise distribution.
@@ VI Conclusion
In this work, we investigate a new task called image denoising adversarial attack,aiming to explore whether deep image denoising methods are robustness to small adversarial perturbations. Based on this idea, we propose the denoising-PGD image denoising adversarial attack and the L2-denoising-PGD attack with constraints to the noise distribution. Through extensive experiments, we verify that the proposed method has consistent excellent adversarial effect as well as transferability among different types of deep image denoising models. Deep image denoising methods are not robust, especially blind denoising models. The model-data hybrid-driven models have shown better robustness comparing to pure data-driven non-blind denoising methods. Unlike adversarial attack tasks targeting image classification, denosing-PGD shows surprising transferability in image denoising. We conjecture and verify that this is due to the high similitude of deep image denoising methods, and the adversarial region can form a continuous space. In addition, we find that the classical model-driven image denoising approach shows resistance to the adversarial attack.

### v2

@@ IV-E Color image denoising adversarial attack
Under the constraint of denoising-PGD, 3-channel color RGB adversarial samples are generated for non-blind denoising model FFDNet with noise level set to 25, maximum perturbation $\varepsilon$ set to 0.055, and 5 iterative steps.
Figure 9: Effect of denoising-PGD on color RGB images (butterfly).
@@
As can be seen in Fig. 9, the addition of adversarial noise does not make a perceptible difference to the noise image, but makes the model denoised substantially less effective. We can see that PSNR decreases by 0.5374 dB, MAE increases by 0.3679, and SSIM decreases by about 0.0139. Visually, the enlarged areas in the denoised image produce the wrong texture, and many image details are erased.
@@
In the experiments to test the transferability, we selected multiple types of deep image denoising models for color RGB image denoising: non-blind denoising methods FFDNet and BRDNet, blind denoising methods Noise2Noise and FAN, plug-and-play methods CurvPnP and DPIR. The adversarial samples generated for FFDNet under denoising-PGD are tested on the above models.
TABLE VII: Transferability of RGB adversarial samples generated by FFDNet to various models on Set3c
Gauss
Adv
Gauss
Adv
Gauss
Adv
Image
PSNR
PSNR
MAE
MAE
SSIM
SSIM
FFDNet
31.06
30.13
5.09
6.00
0.934
0.924
BRDNet
31.22
30.63
5.29
5.73
0.938
0.929
Noise2Noise
29.24
28.87
6.46
6.78
0.913
0.909
FAN
31.58
31.24
4.82
4.89
0.942
0.938
CurvPnP
31.68
30.51
5.02
5.84
0.944
0.915
DPIR
31.60
30.49
5.05
5.90
0.942
0.911
@@
As can be seen from Table VII, the transferability of denoising-PGD on RGB images generated for FFDNet is less than the transferability on grayscale images generated for DnCNN. The adversarial examples show a high adversarial effect on the same type of denoising model, with PSNR decreasing by 0.59 dB, MAE increasing by 0.44, and SSIM decreasing by 0.01 for FFDNet. The adversarial samples show a higher adversarial effect on the plug-and-play model, with PSNR decreasing about 1.1 dB. While for blind denoising model, the adversarial effect decreases, but are still have about 0.3 dB in PSNR. Based on this, we can say that the best transferable adversarial samples can be obtained by applying the image denoising adversarial attack method to a non-blind deep image denoising model. The observation also coincides with the result in section III.C, i.e., the non-blind denoising model has the smallest adversarial space.
IV-F Deep image denoising adversarial training
@@
We used denoisng-PGD at an iteration number of 5 and $\varepsilon=0.012$ to generate the adversarial noise images. Mix the adversarial noise images and Gaussian noise images in 1:1 ratio and feed into the model training phase. The adversarial samples re-generated by denoising-PGD on DnCNN and DnCNN-B are integrated as a test set for testing the resistance of the model after adversarial training against adversarial samples. As can be seen in Fig. 10, the PSNR decremental are reduced from the original 0.8862 dB to 0.1609 dB, and visually denoised adversarial samples become closer to original images. Evaluation of the effect of adversarial training is showed in Table VIII. The difference in the average denoising effect before and after adversarial training is reduced from 0.8 dB to an impressive 0.1 dB. Surprisingly, the adversarial trained denoising model doesn’t show any performance decreasing on original Gaussian noise images, as the main drawback of adversarial training is performance decreasing on the original task. Fig. 11 shows that the adversarial trained model has even better denoising effect on pure Gaussian noise images than vanilla deep image denoising model. The adversarial trained model can reduce artifacts which are commonly appeared in deep image denoising models.
Figure 11: Denoising effect of the adversarial training model and the original model when facing the same Gaussian noise image.
TABLE VIII: Evaluation of the adversarial training.
Cameraman House Pepper Fishstar Monarch Airplane Parrot Lena Barbara Ship Man Couple Avg.(RES)
PSNR 29.8 33.2 30.6 29.2 30.0 28.8 29.0 30.7 28.3 28.7 28.6 28.5 29.6
Gauss SSIM 0.86 0.86 0.88 0.86 0.91 0.86 0.85 0.88 0.82 0.82 0.79 0.82 0.85
MAE 5.2 3.9 5.3 6.4 5.4 5.9 6.1 5.1 7.0 6.6 6.9 6.9 5.9
PSNR 29.9 33.0 30.6 29.3 30.1 29.0 29.3 30.8 28.6 28.9 28.8 28.6 29.7(+0.1)
Adv SSIM 0.87 0.86 0.87 0.87 0.91 0.87 0.85 0.88 0.83 0.82 0.81 0.83 0.86(+0.01)
MAE 5.2 4.0 5.3 6.4 5.4 5.9 6.0 5.2 6.8 6.5 6.8 6.9 5.9(-0.0)
@@ V-A Image denoising adversarial perturbation on clean images
Figure 12: Demonstration of the effect of generating adversarial samples based on the clean and noise images (a) based on the clean image (b) based on the noise image.
@@
In summary, the image denoising network do not have adversarial phenomenon in the case that generating adversarial instances on the clean image. We conjecture that similar to the classification boundary of image classification task, image denoising also has a denoising space, and the outide of the space is the part of the network that is difficult to denoise. Therefore, in this paper, we use the noise map with Gaussian noise added to the input to generate the adversarial samples.
V-B Image denoising adversarial attack on the classic method
@@
Figure 13: Effectiveness of BM3D under image denoising adversarial attacks, where the adversarial samples are generated by DnCNN under denoising-PGD. The test values are PSNR/MAE/SSIM from left to right, respectively.
@@ V-C L2-denoising-PGD image denoising adversarial attack
  INPUT: Gaussian noise sample $x$; its clean image $y$; denoiser $D$; loss function $J$; perturbation size $\varepsilon$; maximum iterations $T$.
  OUTPUT: adversarial sample ${x}^{\prime}$.
@@
V-D The effect of L2-denoising-PGD adversarial attack
In order to eliminate the disturbing pitfall that the random perturbation of the PGD method may have an impact on the noise distribution, the L2-denoising-PGD image denoising adversarial attack method is proposed. As can be seen from the distribution comparison plot in Fig. 14, the adversarial noise of denoising-PGD still has some influence on the distribution, which overlaps more with the Gaussian distribution curve after the L2 constraint is applied. In other words, the noise distribution of denoising-PGD is close to the Gaussian distribution assumption, while the noise distribution of the L2-denoising-PGD is more consistent with the Gaussian distribution assumption after the constraint is applied.
Figure 14: Visual comparison of the noise distribution of the adversarial samples generated under denoising-PGD with the Gaussian noise corresponding to the same noise level. (Upper) denoising-PGD, W-distance=0.0024. (Lower) L2-denoising-PGD, W-distance=0.0017.
Figure 15: The effect diagram of the image denoising adversarial attack (a) denoising-PGD. (b) L2-denoising-PGD.
The L2-denoising-PGD image denoising adversarial attack method also attaches an L2 constraint to the perturbation range of the noise, which further enhances the imperceptibility of the adversarial perturbation and further reduces the impact on the original image. As can be seen from Fig. 15, the effect of the adversarial under such a constraint that conforms to the Gaussian distribution noise assumption and limits the noise perturbation range is very little affected, and the PSNR residual is 0.8051 dB, which is 0.2584 dB lower than the case without the constraint, and the image still has a significant adversarial effect.
In addition, by using the L2-denoising-PGD on different kinds of deep neural denoising networks to generate grayscale adversarial samples and color adversarial samples and testing their adversarial and transferability on the remaining classes of deep neural denoising networks, we can witness that the method is still superior in maintaining the noise distribution, eliminating the concern that the adversarial effect is significantly related to the change of the noise distribution.
@@ VI Conclusion
In this work, we investigate a new task called image denoising adversarial attack, aiming to explore whether deep image denoising methods are robustness to small adversarial perturbations. Based on this idea, we propose the denoising-PGD image denoising adversarial attack and the L2-denoising-PGD attack with constraints to the noise distribution. Through extensive experiments, we verify that the proposed method has consistent excellent adversarial effect as well as transferability among different types of deep image denoising models. Deep image denoising methods are not robust, especially blind denoising models. The pure data-driven non-blind denoising model have shown better robustness comparing to hybrid-driven models and non-blind denoising models. Unlike adversarial attack tasks targeting image classification, denoising-PGD shows surprising transferability in image denoising. We conjecture and verify that this is due to the high similitude of deep image denoising methods, and the adversarial region can form a continuous space. In addition, we find that the classical model-driven image denoising approach shows resistance to the adversarial attack.